Box 3422 c/s  K. Scott Rowe  1.505.835.3637

Socorro, NM 87801 kscott@mailhost.nmt.edu

Job Objective: System Administrator Work Experience

• Manager of Systems Programming, Tech Computer Center, NMIMT August 1998 – Present

• Manage group of systems programmers

• Project manager for programming group

• Implement long-range plans for Computer Center

• System Programmer, Tech Computer Center, NMIMT November 1995 – August 1998

• Install and upgrade UNIX and Windows operating systems.

• Install software for UNIX, NOVELL, Windows systems.

• Design new machine specifications for purchase and support.

• Maintain the WWW server for NMIMT using Apache.

• Ensure uptime of 200+ hosts and servers in a heterogeneous network.

• Responsible for daily backup of 50+ Gig network filesystem using Legato Networker.

• User Consultant, Tech Computer Center, NMIMT May 1994 – November 1995

• Assisted users with UNIX, NOVELL/Windows and MacOS applications.

• Specialized in UNIX, X11, Interleaf, and Emacs.

• Performed troubleshooting on UNIX, NOVELL, MacOS systems.

• System Administrator, Digitized Astronomy, Socorro, NM 87801 December 1993 – May 1994

• Maintained networked Intel machines running UNIX/RT.

• Ported software for the Digitized Astronomy Super Nova Search.

• Computer Science Tutor, ASAP Tutoring, NMIMT May 1991 – December 1993

• Assisted university students in computer science courses.

• Specialized in C, Pascal, and FORTRAN

• User Consultant, Technical Communications Department, NMIMT January 1991 – May 1991

• Assisted users with DOS and MacOS systems and applications.

• Performed troubleshooting on DOS and MacOS systems.

Experience

• Improved and maintain in-house program for installing SunOS, Solaris, Linux, Win95, with associated software packages.

• Designed and implemented 100Mb network of 100+ hosts.

• I have experience administrating SunOS, Linux, NOVELL, OSF/1, Solaris, and Win95. I can program in C, Perl, and Borne Shell.

• Software Construction, Designed and developed web tracking software in a zero defect environment.

• Compilers, Designed and implemented an ICON compiler on SPARC platform.

• Operating Systems, Implemented light-weight processes in the Linux 2.0 kernel, and coroutines in DOS 6.22.

• Computer Architecture, Designed and simulated a 3-stage pipelined processor in Verilog using VHDL.

Education

• Bachelor of Science in Computer Science, New Mexico Institute of Mining and Technology

• Graduation Date August 1997.